Overview

Question 1--/1Question 2--/1Question 3--/1Question 4--/1

Send to Gradebook


Prev

Next

Question 1

--/1

View Policies

Current Attempt in Progress

The ledger of Blue Spruce Corp. on March 31 of the current year includes the selected accounts below before adjusting entries have been prepared.

DebitCredit

Supplies

$7,500

Prepaid Insurance

9,000

Equipment

62,500

Accumulated Depreciation—Equipment

$21,000

Notes Payable

50,000

Unearned Rent Revenue

31,000

Rent Revenue

150,000

Interest Expense

0

Salaries and Wages Expense

35,000

An analysis of the accounts shows the following.

1.The equipment depreciates $700 per month.2.Half of the unearned rent revenue was earned during the quarter.3.Interest of $1,000 is accrued on the notes payable.4.Supplies on hand total $2,125.5.Insurance expires at the rate of $1,000 per month.

Prepare the adjusting entries at March 31, assuming that adjusting entries are made quarterly. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when the amount is entered. Do not indent manually.)

No.

Date

Account Titles and Explanation

Debit

Credit

1.

Mar. 31

enter an account title to record the first transaction

Enter a debit amount

Enter a credit amount

enter an account title to record the first transaction

Enter a debit amount

Enter a credit amount

2.

Mar. 31

enter an account title to record the second transaction

Enter a debit amount

Enter a credit amount

enter an account title to record the second transaction

Enter a debit amount

Enter a credit amount

3.

Mar. 31

enter an account title to record the third transaction

Enter a debit amount

Enter a credit amount

enter an account title to record the third transaction

Enter a debit amount

Enter a credit amount

4.

Mar. 31

enter an account title to record the fourth transaction

Enter a debit amount

Enter a credit amount

enter an account title to record the fourth transaction

Enter a debit amount

Enter a credit amount

5.

Mar. 31

enter an account title to record the fifth transaction

Enter a debit amount

Enter a credit amount

enter an account title to record the fifth transaction

Enter a debit amount

Enter a credit amount

eTextbook and Media

List of Accounts

Save for Later

 

 

Attempts: 0 of 5 used

 

Submit Answer

 

Send to Gradebook


Prev

Next
Answer #1
Account Titles and Explanation Debit Credit
Mar 31 Depreciation expense 2100 =700*3
       Accumulated Depreciation—Equipment 2100
Mar 31 Unearned Rent Revenue 15500 =31000/2
      Rent Revenue 15500
Mar 31 Interest Expense 1000
      Interest Payable 1000
Mar 31 Supplies expense 5375 =7500-2125
     Supplies 5375
Mar 31 Insurance expense 3000 =1000*3
      Prepaid Insurance 3000